Suppose a geyser has a mean time between eruptions of

64 minutes.

Let the interval of time between the eruptions be normally distributed with standard deviation

12 minutes.

Complete parts ​(a) through ​(e) below make sure you awnser A part B part C part D part E awnsers all parts correctly9:47 PM Suppose a geyser has a mean time between eruptions of 64 minutes. Let the interval of time between the eruptions be n
Answer #1

solution: l=64 T = 12 a) PIX>70) = P12) 70-64) p(230.5) 0130857 7***** b) plis 7o) = P(23 70-64 ) = P(2315) 1219 X-M 2 h = x